# GuardSplat: Efficient and Robust Watermarking for 3D Gaussian Splatting

### v2

@@ 4.4 Full Objective
For optimization, we first freeze CLIP’s visual and textual encoders and train the message decoder by minimizing the message loss $\mathcal{L}_{\text{msg}}$ in Eq. 7 between the input and extracted messages. After pre-training the CLIP-guided message decoder, we then employ it to watermark the pre-trained 3DGS models. We freeze the message decoder and utilize it to extract the message from the rendered views, and the secret message can be embedded into 3DGS models by minimizing the following loss as:
@@
Table 2: Comparisons of the start-of-the-art methods on Blender [27] and LLFF [26] datasets for bit accuracy w.r.t various distortion types. We show the results on 16-bit messages. Bold and underline texts indicate the best and second-best performance.
Methods No distortion Noise Rotation Scaling Blur Crop Brightness JPEG Combined
($\mu$=0.1) ($\pm$$\pi$/6) ($\leq$25%) ($\sigma$=$0.1$) (40%) (0.5$\sim$1.5) (10% quality)
@@
WateRF [13] + TensoRF [6] 95.67 95.36 93.13 93.29 95.25 95.40 90.91 86.99 84.12
WateRF [13] + 3DGS [17] 92.07 87.35 88.28 94.33 92.92 95.07 88.71 88.49 86.37
GuardSplat (Ours) + HiDDeN [60] 98.75 96.63 90.02 95.93 94.87 97.25 94.97 90.04 88.70
GuardSplat (Ours) 99.64 99.60 94.56 98.75 99.27 98.71 97.46 94.70 93.38
Figure 6: Visual comparisons between WateRF [13], WateRF + 3DGS, GuardSplat + HiDDeN [60], and our GuardSplat on $N_{L}=32$ bits. Heatmaps show the differences (×$\times 10$) between the watermarked and original views. Red text indicates the best performance.
@@ 5 Experiments
Datasets. Following the experimental settings in [24, 13], we choose 2 commonly-used datasets: the Blender [27] and LLFF [26] datasets, for evaluation. Specifically, the Blender dataset consists of 8 synthetic bounded scenes, while the LLFF dataset consists of handheld forward-facing captures of 8 real scenes. For each scene, we employ 200 test views to evaluate the visual quality and bit accuracy. We report average values across all testing views in our experiments.
Baselines. We compare GuardSplat with 5 baselines to guarantee a fair comparison: 1) CopyRNeRF [24], 2) WateRF [13], 3) HiDDeN [60] + 3DGS [17], 4) WateRF + 3DGS, and 5) GuardSplat + HiDDeN. Specifically, CopyRNeRF and WateRF are NeRF [27] watermarking methods. HiDDeN + 3DGS denotes training 3DGS models on the images watermarked by HiDDeN. WateRF + 3DGS represents applying WateRF to 3DGS models. GuardSplat + HiDDeN indicates replacing our message decoder with HiDDeN’s.
@@
Implementation Details. GuardSplat is implemented on the top of a Pytorch [31] implementation of 3DGS [17] [^1] https://github.com/graphdeco-inria/gaussian-splatting. Our experiments run on a single RTX 3090 GPU with 24G memory. We first train a 3DGS model on a given scene with multi-view photos, and then embed the message into that pre-trained 3DGS model. The size of learnable SH offsets is equal to that of the SH coefficients in the pre-trained models. Our decoder is a multi-layer perceptron (MLP) with 3 fully-connected (FC) layers, where the channels of the first 2 FC layers are separately set to 512 and 256, while the last FC channel is equal to $N_{L}$. For training the message decoder, we employ Adam [18] as the optimizer with a weight decay of $10^{-6}$ and a batch size of $64$. The maximum epoch is set to 100, and the learning rate is set to $5\times 10^{-3}$ as default. It only takes 5 minutes for optimization. For watermarking 3DGS models, we employ Adam [18] as the optimizer with a weight decay of $10^{-6}$ and a batch size of $16$. The maximum epoch is set to 100, and the learning rate of the learnable SH offsets is set to $5\times 10^{-3}$ as default. The hyper-parameters in Eq. 10 are set as $\lambda_{\text{recon}}\!=\!1$, $\lambda_{\text{msg}}\!=\!0.03$, and $\lambda_{\text{off}}\!=\!10$, respectively. Each 3DGS model takes 20 minutes for watermarking.
Evaluation Methods. We follow the standard of digital watermarking to evaluate GuardSplat and the baselines in five aspects: 1) Capacity: We evaluate the bit accuracy across various message lengths $N_{L}\in\{16,32,48\}$ on the 2D rendered views. 2) Invisibility: We evaluate the visual similarity of views rendered from the watermarked and original models using Peak Signal-to-Noise Ratio (PSNR), Structured Similarity Index (SSIM) [48], and Learned Perceptual Image Patch Similarity (LPIPS) [58]. 3) Robustness: We investigate the extraction accuracy with $N_{L}=16$ bits on various visual distortions, including Gaussian Noise, Random Rotation, Random Scaling, Gaussian Blur, Center Crop, Brightness Jittor, 10% JPEG Compression, and a combination of Crop, Brightness, and JPEG. 4) Security: We perform StegExpose [5], an LSB steganography detection, on the rendered views. The detection set is built by mixing the views rendered from the watermarked and original models with equal proportions. 5) Efficiency: We analyze the relationship between bit accuracy and training time.
@@
Figure 7: Performance over different hyper-parameter values with $N_{L}=32$ bits on the Blender [27] and LLFF [26] datasets.
@@ 5.1 Experimental Results
Capacity & Invisibility. We report the capacity and invisibility across various message lengths $N_{L}\in\{16,32,48\}$ on the Blender [27] and LLFF [26] datasets in terms of bit accuracy, PSNR, SSIM, and LPIPS in Table 1. As demonstrated, our GuardSplat surpasses all the competitors with a consistently superior performance w.r.t various message lengths, achieving a significant improvement in extraction accuracy and reconstruction quality. Moreover, GuardSplat achieves much higher bit accuracy than GuardSplat + HiDDeN on $N_{L}\geq 32$ bits, indicating that our CLIP-guided message decoder is superior to HiDDeN’s in handling large-capacity messages. Besides, GuardSplat + HiDDeN outperforms WateRF + 3DGS as they they share the same message decoder, which demonstrates the superiority of our SH-aware watermark embedding module.
Robustness. We report bit accuracy across various visual distortions on $N_{L}=16$ bits in Table 2. Visually, our GuardSplat outperforms all the baselines across various types of distortions, demonstrating that our Anti-distortion Message Extraction module enables the watermarked 3DGS models to learn robust SH features during optimization.
Security. We claim that our GuardSplat is secure. Since it adaptively embeds messages by slightly perturbing the SH features of every 3D Gaussians, it is difficult to remove the watermark from the model file. To further verify the security, we employ StegExpose [5] to detect steganographic content in rendered views. Figure 5 depicts the detection results on our GuardSplat and the baselines. As shown, our GuardSplat achieves superior security to the competitors.
Efficiency. As shown in Figure 2, the efficiency of existing advances is unsatisfactory. Specifically, training a CopyRNeRF requires 84 hours, while WateRF, and WateRF + 3DGS separately spend 14 and 13 hours for optimization as it takes 12 hours to train the HiDDeN. Compared to these methods, our GuardSplat achieves much higher efficiency, which only takes 5 and 20 minutes to train the message decoder and watermark a 3DGS asset, respectively.
Visual Comparison. We visually compare our GuardSplat with baselines on $N_{L}=32$ bits in Figure 6. As shown, our results present superior reconstruction quality and bit accuracy to the competitors. Moreover, the fidelity of WateRF + 3DGS is much lower than GuardSplat + HiDDeN, which demonstrates that altering all the attributes during optimization may lead to a significant decline in image quality.
5.2 Ablation Study & Sensitivity Analysis
We first conduct ablation experiments to prove the effectiveness of our model designs, including various watermark embedding strategies and different loss combinations. Subsequently, we evaluate the bit accuracy and fidelity of our method across different values of $\lambda_{\text{recon}}$, $\lambda_{\text{msg}}$, and $\lambda_{\text{off}}$ to analyze the sensitivity. The results are evaluated on the Blender [27] and LLFF [26] datasets with $N_{L}=32$ bits.
Table 3: Comparisons across various watermark embedding strategies w.r.t $N_{L}=32$ bits on Blender [27] and LLFF [26] datasets. Bold text indicates the best and second-best performance.
@@ 5.2 Ablation Study & Sensitivity Analysis
Table 4: Comparisons across different loss combinations w.r.t $N_{L}=32$ bits on Blender [27] and LLFF [26] datasets. The first row denotes the original 3DGS model, and $\mathcal{L}_{\text{recon}}=\mathcal{L}_{\text{rgb}}$ + $\mathcal{L}_{\text{lpips}}$ indicates the reconstruction loss.
@@ 6 Conclusion
In this paper, we propose GuardSplat, a novel framework to effectively protect the copyright of 3DGS assets. Specifically, we build an efficient message decoder via CLIP-guided Message Decoupling Optimization, enabling the watermarked 3DGS models to achieve high capacity and efficiency. Moreover, we propose a SH-aware message embedding module tailored for 3DGS, which can seamlessly embed the messages while maintaining fidelity, meeting the demands for invisibility and security. We further propose an anti-distortion message extraction module to achieve strong robustness against various visual distortions. Experimental results demonstrate that our GuardSplat outperforms the baselines and achieves fast training speed.
Acknowledgement
This project is supported by the Project of Guangdong Provincial Key Laboratory of Information Security Technology (Grant No. 2023B1212060026), the National Natural Science Foundation of China (12326618), the National Natural Science Foundation of China (U22A2095), and the Key-Area Research and Development Program of Guangzhou (202206030003).
@@ A Overview
  • Section B presents the superior performance of our GuardSplat to the state-of-the-art methods.
  • Section C supplements more implementation details, including the architecture of the message decoder and distortion layers.
@@
  • Section D conducts an additional evaluation for security, exploring whether the watermarks can be simply removed from model files.
  • Section E visualizes more experimental results, including visual comparisons of various ablations in Tables 3 and 4 of the main paper, and quantitative results on larger-capacity messages $N_{L}=\{64,72\}$.
B Superiority of our GuardSplat
To prove the superiority of our GuardSplat, we show the performance between our GuardSplat and state-of-the-art methods in Figure S1, where our GuardSplat + HiDDeN [60] denotes replacing our CLIP-guided message decoder with HiDDeN’s to optimize our GuardSplat. As shown, our GuardSplat and GuardSplat + HiDDeN achieve superior performance in bit accuracy and reconstruction quality to the compared methods, demonstrating that the proposed watermarking watermark can seamlessly embed the messages into 3DGS assets while maintaining the fidelity. Moreover, unlike the compared methods that require hours and even days for optimization, the training time of our GuardSplat is 25 minutes, which is more suitable for real-world applications.
C More Implementation Details
We supplement more details of our GuardSplat.
Distortion layers. Since the proposed CLIP-guided message decoder can deal with Gaussian Noise and Gaussian Blur, it only considers the visual distortions of cropping, scaling, rotation, brightness jittering, and JPEG compression. The former four visual distortions are differentiably re-implemented by Pytorch [31] built-in functions, while the differentiable JPEG comparison is built by [37]. During training, we employ the differentiable distortion layer to simulate various visual distortions for optimization. In the test, the rendered views are distorted using the visual distortions built by OpenCV for evaluation.
Figure S1: Bit accuracy and PSNR of CopyRNeRF [24], WateRF [13], WateRF [13] + 3DGS [17], our GuardSplat + HiDDeN [60] and our GuardSplat with $N_{L}=32$ bits on Blender [27] and LLFF [26] datasets. The radius of circles is proportional to their total training time evaluated on RTX 3090 GPU.
@@ C More Implementation Details
Figure S2: The architecture of our message decoder. Given the output features $F_{\mathcal{T}}$ or $F_{\mathcal{V}}$, we first pass it through 2 FC layers with GELU activations, where the first and second FC layers have 512 and 256 channels, respectively. Then, we map the output feature to the binary message by using a $N_{L}$-channel FC layer and a Sigmoid activation.
Architecture of Message Decoder. Figure S2 depicts the architecture of our CLIP-guided Message Decoder. Thanks to CLIP’s rich representation, our decoder can achieve excellent performance with minimal parameter size.

### v5

@@ 4.4 Full Objective
For optimization, we first freeze CLIP’s visual and textual encoders and train the message decoder by minimizing the message loss $\mathcal{L}_{\text{msg}}$ in Eq. 7 between the input and extracted messages. After pre-training the CLIP-guided message decoder, we then employ it to watermark the pre-trained 3DGS models. We freeze the message decoder and utilize it to extract the message from the rendered views, and the secret message can be embedded into 3DGS models by minimizing the following loss:
@@
Table 2: Comparisons of the start-of-the-art methods on Blender [32] and LLFF [31] datasets for bit accuracy w.r.t various distortion types. We show the results on 16-bit messages. Bold text indicates the best performance in this table.
Methods None Noise Rotation Scaling Blur Crop Brightness JPEG VAE Attack [68] Combined
($\mu$=0.1) ($\pm$$\pi$/6) ($\leq$25%) ($\sigma$=$0.1$) (40%) (0.5$\sim$1.5) (10% quality) Bmshj2018
@@
WateRF [14] 95.67 95.36 93.13 93.29 95.25 95.40 90.91 86.99 51.73 84.12
3DGS [19] w/ WateRF [14] 92.89 87.35 88.28 90.33 91.92 89.07 88.71 88.49 55.48 86.37
GaussianMarker [13] 99.36 99.13 70.84 97.89 94.40 98.52 95.78 86.22 52.00 83.49
GuardSplat (Ours) + CIN [29] 95.75 94.87 90.89 94.50 95.16 93.82 93.97 88.61 49.25 84.03
GuardSplat (Ours) + SSL [9] 99.50 99.57 86.78 84.53 98.79 77.54 94.31 92.99 47.42 74.85
GuardSplat (Ours) + HiDDeN [69] 98.75 96.63 90.02 95.93 94.87 97.25 94.97 90.04 53.14 88.70
GuardSplat (Ours) + StegaStamp [52] 99.00 98.38 53.21 95.17 98.17 51.34 95.48 88.81 80.12 64.75
GuardSplat (Ours) 99.64 99.60 94.56 98.75 99.27 98.71 97.46 94.70 82.35 93.38
@@ 5 Experiments
Datasets. Following the settings in [28, 14], we choose two commonly-used datasets: the Blender [32] and LLFF [31] datasets, for evaluation. Specifically, the Blender dataset consists of 8 synthetic bounded scenes, while the LLFF dataset consists of handheld forward-facing captures of 8 real scenes. For each scene, we employ 200 test views to evaluate the visual quality and bit accuracy. We report average values across all testing views in our experiments.
Baselines. We compare GuardSplat against six baselines to ensure a fair comparison: 1) CopyRNeRF [28], 2) WateRF [14], 3) GaussianMarker [13], 4) 3DGS w/ WateRF, 5) 3DGS trained on watermarked images (i.e., 3DGS + 2D watermarking methods), and 6) GuardSplat optimized by other pre-trained 2D watermarking decoders, (i.e., GuardSplat + 2D watermarking decoders). Specifically, CopyRNeRF and WateRF are NeRF-based [32] watermarking methods, while GaussianMarker is a 3DGS watermarking approach. 3DGS w/ WateRF refers to applying WateRF to 3DGS models. The 2D watermarking methods include DwtDctSvd [34], StegaStamp [52], SSL [9], and CIN [29].
Figure 5: Visual comparisons with $N_{L}=32$ bits. Heatmaps show the differences (×$\times 10$) between the watermarked and original views.
@@
Implementation Details. GuardSplat is implemented on the top of a Pytorch [37] implementation of 3DGS [19] [^1] https://github.com/graphdeco-inria/gaussian-splatting. Our experiments run on a single RTX 3090 GPU. We first train a 3DGS model on a given scene with multi-view photos and then embed the message into that pre-trained 3DGS model. The size of learnable SH offsets is equal to that of the SH coefficients in the pre-trained models.
For training the message decoder, we employ Adam [20] as the optimizer with a weight decay of $10^{-6}$ and a batch size of $64$. The maximum epoch is set to 100, and the learning rate is set to $5$×$\times 10^{-3}$ as default. Given the message length $N_{L}$, we randomly select $\mathtt{min}(2^{N_{L}},N_{K})$ samples from a total of $2^{N_{L}}$ messages as training and test data, where $N_{K}$=$2048$. It only takes 5 minutes for optimization. Thanks to CLIP’s rich representation, our decoder achieves excellent performance with only 3 FC layers, detailed architecture is provided in Supp. B.
For watermarking 3DGS models, we employ Adam [20] as the optimizer with a weight decay of $10^{-6}$ and a batch size of $16$. The maximum epoch is set to 100, and the learning rate of the SH offsets is set to $5$$\times$$10^{-3}$. The hyper-parameters in Eq. 10 are set as $\lambda_{\text{recon}}$=$1$, $\lambda_{\text{msg}}$=$0.03$, and $\lambda_{\text{off}}$=$10$, respectively. It takes 10 minutes for watermarking.
For Distortion Layer, since the proposed CLIP-guided message decoder can deal with Gaussian Noise and Gaussian Blur, we only consider the visual distortions of cropping, scaling, rotation, brightness jittering, JPEG compression, and VAE attacks. The former four visual distortions are differentiably re-implemented by Pytorch [37] built-in functions, while the differentiable JPEG comparison is built by [43]. During training, we employ the differentiable distortion layer to simulate various visual distortions for optimization. In the test, the rendered views are distorted using the visual distortions built by OpenCV for evaluation.
@@
Figure 6: Performance across different hyper-parameter values with $N_{L}=32$ bits on the Blender [32] and LLFF [31] datasets.
Evaluation Metrics. We follow the standard of digital watermarking to evaluate GuardSplat and the baselines in five aspects: 1) Capacity: We evaluate the bit accuracy across various message lengths $N_{L}\in\{16,32,48\}$ on the 2D rendered views. 2) Invisibility: We evaluate the visual similarity of views rendered from the watermarked and original models using Peak Signal-to-Noise Ratio (PSNR), Structured Similarity Index (SSIM) [56], and Learned Perceptual Image Patch Similarity (LPIPS) [65]. 3) Robustness: We investigate the extraction accuracy with $N_{L}=16$ bits across various visual distortions, including Gaussian Noise of $\mu$=0 and $\sigma$=0.1, Random Rotation of angles within $[-\frac{\pi}{6},+\frac{\pi}{6}]$, Random Scaling of ratios within $[0.75,1.25]$, Gaussian Blur of kernel_size=3 and $\sigma$=0.1, 40% Center Crop, Brightness Jittoring of ratios within $[0.5,1.5]$, JPEG compression of 10% image quality, VAE attack [68] based on Bmshj2018, and the combination of Crop, Brightness, and JPEG. 4) Security: We perform StegExpose [6], an LSB steganography detection, on the rendered views. The detection set is built by mixing the watermarked and original views with equal proportions. 5) Efficiency: We analyze the relationship between bit accuracy and training time.
@@ 5.1 Experimental Results
Security. We claim that our GuardSplat is secure. Since it adaptively embeds messages by slightly perturbing the SH features of every 3D Gaussian, it is difficult to remove the watermark from the model file. To further verify the security, we employ StegExpose [6] to detect steganographic content in rendered views. Figure 4 depicts the detection results on our GuardSplat and the baselines. As shown, our GuardSplat achieves superior security to the competitors.
Capacity & Invisibility. We report the capacity and invisibility across various message lengths $N_{L}$$\in$$\{16,32,48\}$ on the Blender [32] and LLFF [31] datasets in terms of bit accuracy, PSNR, SSIM, and LPIPS in Table 1. As demonstrated, our GuardSplat surpasses all the competitors with a consistently superior performance w.r.t various message lengths, achieving a significant improvement in extraction accuracy and invisibility. Moreover, GuardSplat achieves higher bit accuracy than GuardSplat + 2D watermarking decoders with $N_{L}\geq 32$ bits, indicating that our CLIP-guided message decoder is better at handling large-capacity messages. Besides, GuardSplat + HiDDeN outperforms 3DGS w/ WateRF as they share the same decoder, proving the superiority of our SH-aware Message embedding module.
Robustness. We report bit accuracy across various visual distortions with $N_{L}=16$ bits in Table 2. Visually, our GuardSplat outperforms all the baselines across various distortions, demonstrating that our Anti-distortion Message Extraction module enables the watermarked 3DGS models to learn robust SH features during optimization.
Efficiency. As shown in Figure 2, the efficiency of existing advances is unsatisfactory. Specifically, training a CopyRNeRF requires 84 hours, while WateRF, and 3DGS w/ WateRF separately spend 14 and 13 hours for optimization as it takes 12 hours to train the HiDDeN. Compared to these methods, our GuardSplat achieves much higher efficiency, which only takes 5 and 10 minutes to train the message decoder and watermark a 3DGS asset, respectively. We also provide the comparison of watermarking speed in Supp. B.
Table 3: Various embedding methods on Blender [32] and LLFF [31] datasets with $N_{L}$=$32$ bits. Bold text denotes the best score.
@@
Visual Comparison. We visually compare our GuardSplat against baselines with $N_{L}=32$ bits in Figure 5. As shown, our results present superior reconstruction quality and bit accuracy to the competitors. Moreover, the fidelity of 3DGS w/ WateRF is much lower than GuardSplat + HiDDeN, which demonstrates that altering all the attributes during optimization may lead to a significant decline in image quality.
5.2 Ablation Study & Sensitivity Analysis
We first conduct ablation experiments to prove the effectiveness of our model designs, including various message embedding strategies and loss combinations. Subsequently, we evaluate the bit accuracy and fidelity of our method across different values of $\lambda_{\text{recon}}$, $\lambda_{\text{msg}}$, and $\lambda_{\text{off}}$ to analyze the sensitivity. The results are evaluated on the Blender [32] and LLFF [31] datasets with $N_{L}=32$ bits.
@@ 5.2 Ablation Study & Sensitivity Analysis
Table 4: Different loss combinations with $N_{L}$=$32$ bits on Blender [32] and LLFF [31] datasets. The first row denotes the original 3DGS, and $\mathcal{L}_{\text{recon}}=\mathcal{L}_{\text{rgb}}$ + $\mathcal{L}_{\text{lpips}}$ indicates the reconstruction loss.
@@ 6 Conclusion
In this paper, we present GuardSplat, a novel watermarking framework to protect the copyright of 3DGS assets. Specifically, we build an efficient message decoder via CLIP-guided Message Decoupling Optimization, enabling high-capacity and efficient 3DGS watermarking. Moreover, we tailor a SH-aware message embedding module for 3DGS to seamlessly embed the messages while maintaining fidelity, meeting the demands for invisibility and security. We further propose an anti-distortion message extraction module to achieve strong robustness against various visual distortions. Experiments demonstrate that our GuardSplat outperforms the baselines and achieves fast training speed.
Acknowledgement. This project is partially supported by the NSFC (U22A2095), the Major Key Project of PCL under Grant PCL2024A06, and the Project of Guangdong Provincial Key Laboratory of Information Security Technology (Grant No. 2023B1212060026).
@@ A Overview
  • Section B depicts the architecture of our message decoder guided by CLIP [41], and we also conduct a comparison for watermarking efficiency against the state-of-the-art methods.
@@
  • Section C conducts an additional evaluation for security, exploring whether the watermarks can be simply removed from model files.
  • Section D illustrates the visualization results of various ablations in Tables 3 and 4 of the main paper.
  • Section E reports more results, including the quantitative results on larger-capacity messages $N_{L}=\{64,72\}$, bit accuracy across various rendering situations, and the zoomed-in rendering results between watermarked and original views.
B Decoder Architecture and Watermarking Speed
As shown in Fig. S1, our message decoder only consists of 3 fully-connected (FC) layers, which can accurately map the CLIP textual features to the corresponding binary messages after a 5-minute optimization. Thanks to CLIP’s rich representation, our decoder can achieve excellent performance with minimal parameter size. We also investigate the watermarking efficiency between our GuardSplat and state-of-the-art methods. As shown in the training accuracy curve in Figure S2, our GuardSplat achieves the highest efficiency, which only takes 10 minutes to watermark a pre-trained 3DGS asset.
C Additional Evaluation for Security
We conduct additional experiments to evaluate the security of our GuardSplat in Table S1, investigating whether the malicious users can remove the watermarks from the model file by pruning the $K\%$ of Gaussians, where $K\in\{5,10,15,20,25\}$. “Bottom $K$” denotes pruning $K$ of low-opacity Gaussians, while “random” denotes randomly pruning $K$ of the Gaussians. As demonstrated, our GuardSplat still achieves a bit accuracy of 98.74% when 25% of the low-opacity Gaussians are removed, indicating that simply removing low-opacity Gaussians does not effectively attack our method. Though randomly removing the Gaussians can lead to a significant decline in bit accuracy, it also greatly affects the reconstruction quality (i.e., PSNR, SSIM, and LPIPS), resulting in low-fidelity rendering. This experimental result demonstrates that the malicious cannot directly remove the watermarks from the model file, verifying the security of our GuardSplat.
@@ C Additional Evaluation for Security
Figure S1: The architecture of our message decoder. Given an output feature $F_{\mathcal{T}}$ or $F_{\mathcal{V}}$, we first pass it through two FC layers with GELU activations, where their channels are set to 512 and 256, respectively. Then, we map the feature to the binary message using a $N_{L}$-channel FC layer and a Sigmoid activation.
Figure S2: Training accuracy curves with $N_{L}=32$ bits on Blender [32] dataset. Our GuardSplat achieves high training efficiency, which only takes 10 minutes to watermark a 3D asset.
D Additional Visual Comparisons
D.1 Various Message Embedding Strategies
In the main paper, we explore the performance under various message embedding strategies with $N_{L}=32$ bits (see quantitative results in Table 3). For better comparisons, we further visualize the results of various message embedding strategies in Figure S3. As shown, the proposed SH-aware module achieves superior bit accuracy and reconstruction quality to the competitors.
D.2 Various Loss Combinations
In the main paper, we quantitatively compare the performance across various loss combinations in Table 4. We also conduct a visual comparison of these ablation variants in Figure S4. As shown, “$\mathcal{L}_{\text{\text{recon}}}+\mathcal{L}_{\text{\text{msg}}}+\mathcal{L}_{\text{\text{off}}}$” achieves the best performance in bit accuracy and reconstruction quality.